An investment will pay $150 at the end of each of the next 3 years, $250 at the end of Year 4, $300 at the end of Year 5, and $500 at the end of Year 6. If other investments of equal risk earn 6% annually, what is its present value? Its future value? Do not round intermediate calculations. Round your answers to the nearest cent.

Present value:

Future value:
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

=150/1.06+150/1.06^2+150/1.06^3+250/1.06^4+300/1.06^5+500/1.06^6

=$1175.63(Approx).

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value=1175.63*(1.06)^6

=$1667.66(Approx).
Answer #2

SOLUTION :


Discount rate, r = 6% = 0.06 

=> 1 = r = 1.06


PV 

= 150/1.06+150/1.06^2+150/1.06^3+250/1.06^4+300/1.06^5+500/1.06^6

= 1175.63 ($) (ANSWER).


FV 

= PV*1.06^6 

= 1175.63*1.06^6

= 1667.66 ($) (ANSWER)

answered by: Tulsiram Garg

> 2nd line, please read " 1 = r = 1.06 " as " 1 + r = 1.06 "

Tulsiram Garg Thu, Oct 14, 2021 2:50 AM